A bus is moving at 5.0 m/s when a passenger walks forward at 1.5 m/s relative to the bus. How fast is this passenger moving as seen and measured by (a) a person waiting at the bus stop and (b) the driver of the bus?

1
Expert's answer
2022-05-02T13:43:08-0400

Given:

"v_b=5.0\:\rm m/s"

"v_p=1.5\:\rm m/s"

(a)

"V_r=v_p=1.5\:\rm m/s"

(b)

"V_r=v_b+v_p=5.0+1.5=6.5\:\rm m/s"